Operations summary

The vessel arrived on location at 2230 h on 20 February 2012 after making a rapid passage from Lisbon, Portugal, with an average transit speed of 12.37 kt over the 1713 nmi distance. The R/V JOIDES Resolution’s excellent transit speed can most likely be attributed to a recently cleaned hull, newly polished propellers, and favorable winds and seas. During the transit to the first site, the drilling and logging equipment was inspected and tested to ensure performance during Expedition 340T operations. Prior to the start of logging operations, a presite meeting was held with Siem Offshore staff, IODP staff, the Chief Scientist, and other critical staff.

On arrival at Site U1309, the speed of the vessel was reduced and bridge control was shifted to DP control at 2257 h. Drill floor operations began immediately after switching to DP control. First, the upper guide horn was laid out, the BHA for logging operations was picked up and assembled, and the drilling string was run to the seafloor. Despite preoperational checks, the iron roughneck torquing system failed during the initial pipe trip. While troubleshooting the hydraulic problem, tripping continued using manual tongs and a rope to spin in the drill pipe. The vibration-isolated television (VIT) frame was installed while tripping to the bottom, and at 1015 h on 21 February we began spacing out the bit for reentry into Hole U1309D. What appeared to be the reentry cone was identified almost immediately, and the vessel was positioned for a reentry at ~1115 h. The reentry attempt was made but was unsuccessful, primarily because the assumed reentry cone was in fact a conelike seafloor structure. A bottom survey was then conducted to locate the Hole U1309D reentry cone. It was located and reentered at 1320 h on 21 February. After reentry the logging bit was positioned at 1711 mbrf, which is below known problem sections in Hole U1309D.

The triple combo tool string was rigged up by 1440 h on 21 February and lowered to the hole. The tool string was run into the hole, recording standard measurements as well as an in situ temperature profile from the surface to 3060 mbrf. At 0050 h on 22 February, the tool string was completely back in the pipe but unable to pass through the BHA. After working the logging string for >4 h, the decision was made to use the Kinley crimper/cutter to crimp and cut the logging cable. After deploying two crimping systems and a cutter, the cable was cut and retrieved. The VIT frame was then run to the seabed to verify that the tools were present at the end of the bit. The drill bit cleared the reentry cone at 1410 h. The bottom of the logging tool string was visually observed hanging out of the end of the bit. The drill string was then pulled to surface and a stand of the BHA was set back. The logging tools were removed from the last drill collar with care. All tools were successfully recovered including the radioactive source. It was immediately apparent that both centralizers had failed. The center section of the centralizers appears to have been severely abraded by the formation, likely breaking during pipe reentry. After rigging down the logging tools, the drill string was tripped back into the hole. The VIT frame was then deployed but had to be retrieved for repairs after one of the lights failed. On redeployment the lighting system failed again. The VIT frame was again pulled back and an electronics pod was changed. At 0515 h on 23 February, the VIT frame was again installed and then run to bottom. The vessel was positioned for reentry into Hole U1309D, which occurred at 0705 h. The drill string was run into the hole to 1711 mbrf. The VIT frame was retrieved.

A protected species watch was initiated at 0630 h on 23 February, and the seismic source (a G-gun parallel cluster of two 250 in3 air guns) was deployed 7 mbsl, in preparation for a VSP experiment. After 1 h with no protected species sightings, the seismic guns were soft-started and ramped up to full pressure over the course of the next 30 min. The guns then remained on standby but were fired at least once every 30 min until the VSP experiment began. The VSI tool string was rigged up at 0815 h and run into the hole. The first seismic station was established at 1805 mbrf, and the first shot of the VSP was fired. After a successful first station, the tool string was run to the bottom of the hole to the deepest station at 3016 mbrf. At 3016 mbrf and four successively shallower stations, data indicated that the VSI was not clamping properly. After a period of troubleshooting, the tool was pulled back to surface, where it was confirmed that the caliper was broken. A second VSI sonde was substituted into the tool string, and it was run back into the hole for 1 h of logging before sunset. Three stations were completed, and VSP operations ended for the day because of darkness. The VSI tool string was pulled back to surface and rigged down by 1910 h.

The DSI (sonic) tool string was run next. Given that it used the same centralizers as the triple combo, a decision was made to only log the deep portion of the hole that had not been logged during Expedition 305. The logging bit was tripped down to 2356 mbrf, or 700 mbsf. The pipe trip identified further ledges at 1723, 1726, and 1740 mbrf. The passive heave compensator had to be opened up to pass these ledges. At 2200 h, the DSI tool string was rigged up and run into the hole. The tool string started taking weight inside the BHA and could not be run down to the depth of the drill bit. After unsuccessfully working the tool string for 1 h to pass the BHA, the tool string was pulled back to surface and rigged down at 0330 h on 24 February. Initially there was a concern that something was either damaged or obstructing the BHA. An extended core barrel (XCB) was rigged up and run on wireline to verify that the BHA was free and clear to the bit. After verifying that the BHA was clear, the investigation turned to the DSI tool string. An obstruction was identified at one of the centralizers. The centralizers were cleaned of all rubberized backing material, and this allowed the sonic tool string to clear the landing seat.

While the centralizers were being fixed, the VSI tool string was rigged up and run into the hole to begin the second day of the VSP experiment. The protected species watch began again at 0630 h, and the G-gun cluster was deployed and ramped up following the same method as the previous day. The VSI tool string was run into the hole at 0725 h on 24 February. The VSP experiment continued through the day and was completed prior to dusk. The tool string was then pulled to the surface and rigged down. The logging bit was tripped to 2356 mbrf, or 700 mbsf past the identified ledges at 1723, 1726, and 1740 mbrf. Again, the passive heave compensator had to be opened up to pass these ledges.

At 2020 h, the DSI tool string was rigged up for a second deployment. The tool string was successfully run through the BHA to 3040 mbrf and back up to the surface with no difficulties. The logging run was concluded at 0435 h on 25 February, when the tools were rigged down on the drill floor.

The drill pipe was tripped back to 1759 mbrf so that the upper portion of the hole could be logged with the MSS tool string, which was the final logging run in Hole U1309D. The MSS tool is temperature-limited to 80°C, so the tool string was run into the hole and recorded data from a maximum depth of 2419 mbrf up to the drill pipe. The tool string was run back down, and a repeat pass was recorded from 2165 mbrf up to the surface. By 1030 h, the tools were back on the surface and rigged down. The drill string was pulled clear of the reentry cone at 1055 h and spaced out for a bottom survey.

A seafloor survey was started from Hole U1309D by moving 5 m east, then 5 m south, and then 10 m west in an expanding spiral. The survey was concluded after 1.5 h. An attempt was made to retrieve a sample from a moundlike feature on the seafloor at 30°10.1179′N, 042°07.1118′W. The bit was set down inside the conelike opening of the feature, and a modified advanced piston core (APC) core barrel was lowered by wireline and run into the bottom twice to attempt to sample the material inside the feature. After IODP Hole U1392A was spudded at 1447 h, the core barrel was retrieved and laid out on the drill floor, and the sample was extracted for processing by the technical staff. The drilling string was then pulled back to the drill floor, and the drill collars were secured in the their racks, the upper guide horn was reinstalled, and the rig floor was secured for transit at 2030 h. The thrusters were then raised, and a magnetometer survey was started. The expedition ended with the JOIDES Resolution running two crossing lines with the magnetometer above Site U1309 and preparing to get under way to San Juan, Puerto Rico.